Simplifying the expression 4/9 -1/2 +1/3. Gives?


1
Expert's answer
2021-04-28T05:54:31-0400

"\frac{4}{9}-\frac{1}{2}+\frac{1}{3}=\frac{8}{18}-\frac{9}{18}+\frac{6}{18}=\frac{5}{18}."